### v2

# MCP-Zero: Proactive Toolchain Construction for LLM Agents from Scratch

Abstract
Function-calling has enabled large language models (LLMs) to act as tool-using agents, but injecting thousands of tool schemas into the prompt is costly and error-prone. We introduce MCP-Zero, a proactive agent framework that lets the LLM itself decide when and which external tools to retrieve, thereby assembling a task-specific toolchain from scratch. The framework is built upon three components: (1) Proactive Tool Request, where the model emits a structured $\operatorname{tool\_assistant}\right>$ block that explicitly specifies the desired server and task; (2) Hierarchical Vector Routing, a coarse-to-fine retrieval algorithm that first selects candidate servers and then ranks tools within each server based on the semantic similarity; (3) Iterative Proactive Invocation, enabling multi-round, cross-domain toolchain construction with minimal context overhead, and allowing the model to iteratively revise its request when the returned tools are insufficient. To evaluate our approach we also compile MCP-tools, a retrieval dataset comprising 308 MCP servers and 2,797 tools extracted from the official Model-Context-Protocol repository and normalized into a unified JSON schema. Experiments show that MCP-Zero (i) effectively addresses the context overhead problem of existing methods and accurately selects the correct tool from a pool of nearly 3,000 candidates (248.1k tokens); (ii) reduces token consumption by 98% on the APIbank while maintaining high accuracy; and (iii) supports multi-turn tool invocation with consistent accuracy across rounds. The code and dataset will be released soon at https://github.com/xfey/MCP-Zero.
1 Introduction
The rapid advancement of large language models (LLMs) has catalyzed a paradigm shift from pure text understanding and generation to sophisticated tool-using agents capable of interacting with external systems [30, 8, 1]. With the introduction of function-calling mechanisms, LLMs have transcended the boundaries of their parametric knowledge, enabling them to leverage external tools, third-party APIs, and code execution environments to accomplish complex reasoning chains and real-world tasks [21, 26, 14, 18].
However, as the ecosystem of available tools continues to expand exponentially, the conventional approach of injecting comprehensive tool schemas into the system prompt has emerged as a critical bottleneck (Figure LABEL:fig:teasera). Mainstream practices typically embed complete JSON-Schema of all available tools within the system prompt, leading to substantial context overhead, as shown in Figure 1. For instance, the GitHub MCP server encompasses 26 tools requiring over 4,600 tokens, substantially compressing the available context window for actual task content.
To mitigate context window constraints, recent approaches have adopted retrieval-based strategies that match and inject only the most relevant tool based on semantic similarity with user queries [7, 15]. However, these methods face significant limitations when dealing with complex, multi-step tasks that require coordination across multiple domains (Figure LABEL:fig:teaserb). For example, a query like “Debug the file” requires coordinating tools across different domains: $\operatorname{filesystem}$ for file access, $\operatorname{code\_generation}$ for updates, and $\operatorname{command\_execution}$ for debugging. Single-tool retrieval fails to capture such complex workflows, as the initial query cannot determine all required tools across domains.
In summary, while retrieval methods demonstrate efficacy in single-step, single-domain scenarios, they exhibit fundamental limitations in realistic multi-turn, multi-domain agent environments: (1) passive retrieval - external systems select tools based on initial query rather than allowing models to actively express their evolving needs; (2) semantic misalignment between colloquial user inputs and formal API documentation creates distributional mismatches that reduce retrieval precision; and (3) single-round invocation - tool retrieval is performed only once per query, failing to accommodate the progressive refinement of sub-task requirements or iterative correction when initially retrieved tools prove inadequate.
{mdframed}
[linewidth=1pt, linecolor=gray!50, backgroundcolor=gray!5] <function>
{"description": "Search for GitHub repositories", "name": "mcp_github_search_repositories", "parameters": {"\$schema": "http://json-schema.org/draft-07/schema#", "additionalProperties": false, "properties": {"page": {"description": "Page number for pagination (default: 1)", "type": "number"}, "perPage": {"description": "Number of results per page (default: 30, max: 100)", "type": "number"}, "query": {"description": "Search query (see GitHub search syntax)", "type": "string"}}, "required": ["query"], "type": "object"}}
</function>
Figure 1: Example of a single MCP tool definition from the GitHub MCP server. This tool requires 143 tokens, while the complete server requires over 4,600 tokens.
Modern LLMs possess powerful capabilities in chain-of-thought reasoning, self-reflection, and planning. Rather than having external systems passively select tools based on initial queries, we propose letting the model proactively analyze the context, identify capability gaps, and request tools when external assistance is needed. Based on this insight, we introduce MCP-Zero (Figure LABEL:fig:teaserc), a proactive retrieval framework with the following key components:
Proactive Tool Request. Unlike traditional approaches that passively wait for external retrieval, we return the authority of tool requirement specification to the LLM itself. When external tool assistance is needed, the model proactively generates structured tool request declarations in the following format:
<tool_assistant>
   server: ...  # Platform/permission domain
   tool: ...    # Operation type + target
</tool_assistant>
This mechanism enables spontaneous need expression while ensuring semantic consistency between requests and tool documentation, avoiding colloquial query issues and improving retrieval quality.
Hierarchical Vector Routing. The system employs two-stage retrieval: first filtering candidate servers by platform requirements, then matching specific tools within selected servers. Only top-k tool descriptions are returned, significantly reducing context overhead.
Iterative Proactive Invocation. The model can initiate multiple tool requests throughout the conversation for different subtasks, enabling cross-server toolchain construction from scratch, while introducing only necessary tools at each step. If returned tools are insufficient, the model can refine its requests and reinitiate retrieval, providing fault tolerance and self-correction capabilities.
To support our approach, we also compile MCP-tools, the first retrieval dataset in the MCP domain, comprising 308 servers and 2,797 tools extracted from the official Model-Context-Protocol repository111github.com/modelcontextprotocol/servers and normalized into a unified JSON schema. Our comprehensive experiments demonstrate that MCP-Zero effectively addresses the context overhead problem while maintaining high accuracy in tool selection from nearly 3,000 candidates, reduces token consumption by 98% compared to existing methods, and supports consistent multi-turn tool invocation across conversation rounds.
Our main contributions are summarized as follows:
  • We propose MCP-Zero, enabling LLMs to proactively analyze tasks, iteratively request tools on demand, and dynamically construct multi-step toolchains with minimal context overhead and high accuracy.
  • We design a hierarchical vector routing mechanism that effectively improves tool retrieval accuracy and context utilization efficiency.
  • We construct and release MCP-tools, a comprehensive retrieval dataset covering 308 servers and 2,797 tools from the official Model-Context-Protocol repository.
  • Extensive experiments demonstrate that our method significantly reduces context overhead while maintaining consistent performance, effectively addressing the context length problem in tool-calling scenarios.
2 Related Work
2.1 Tool-Augmented LLMs
The evolution of tool-augmented large language models has progressed through distinct paradigms, each addressing different aspects of the fundamental challenge: how to effectively integrate external capabilities with language model reasoning.
Early Task-Specific Integration. Initial approaches focused on hard-coding specific tools into language models. MRKL [10] introduced modular reasoning by combining neural networks with symbolic modules like calculators, while WebGPT [16] demonstrated web browsing capabilities for information retrieval tasks. Though effective within their domains, these systems suffered from limited scalability and poor generalization to new tool types.
Universal Agent Protocols. The introduction of ReAct [31] marked a paradigm shift by establishing the “observation-action-thought” pattern, creating a universal protocol for tool-augmented reasoning. This framework became the foundation for modern agent systems including LangChain Agents [4] and AutoGen [29], enabling flexible and extensible tool integration architectures.
Training-Based Tool Learning. Parallel developments explored learning tool usage through model training. Toolformer [26] pioneered self-supervised learning for API call generation, teaching models to insert tool invocation markers within natural text generation. Gorilla [18] extended this approach by constructing large-scale instruction datasets with tool calls, enabling supervised learning of query-to-tool mappings across diverse API collections.
Context-Based Tool Injection. The emergence of ChatGPT Function Calling and systems like HuggingGPT [27] introduced context-based approaches, which inject JSON-Schema tool descriptions into system prompts, eliminating the need for specialized training. ART [17] further refined this paradigm by constructing demonstration libraries in chain-of-thought format, leveraging in-context learning for tool selection and usage.
Fundamental Limitations. Despite these advances, existing approaches face critical scalability challenges. Training-based methods require expensive retraining for each toolset update, limiting their adaptability to evolving tool ecosystems. Context-based methods, while more flexible, suffer from prohibitive context overhead when injecting comprehensive tool descriptions—a problem that becomes acute as tool collections scale to thousands of APIs. Most critically, when relevant tools are absent from the predefined context, task completion becomes impossible, highlighting the need for dynamic, on-demand tool discovery mechanisms that can adapt to diverse and evolving task requirements.
2.2 Tool Retrieval for LLMs
The success of Retrieval-Augmented Generation (RAG) in addressing knowledge limitations through the “retrieve-insert-generate” paradigm has inspired its adaptation to tool selection for LLMs [11]. Classical RAG frameworks like REALM [9], RETRO [3] and In-Context RALM [25] demonstrated the effectiveness of dynamically incorporating external knowledge into generation processes. Recent work has extended this paradigm to tool selection, aiming to reduce context overhead by retrieving only the most relevant tools for a given query.
Semantic Similarity-Based Retrieval. Early approaches focused on direct semantic matching between user queries and tool descriptions. Gorilla [18] constructed vector databases from API documentation and usage examples, employing semantic similarity to retrieve relevant tools. Tool2Vec [15] addressed the semantic gap between user requests and formal API descriptions by pre-collecting diverse user invocation patterns and computing averaged embeddings, though this approach requires extensive user interaction datasets for training. RAG-MCP [7] performs server-level matching between user queries and documentation from MCPBench [13], returning all tools from the most similar server as the context of LLMs.
Hierarchical Tool Retrieval. Recognizing the limitations of flat retrieval, several works have explored coarse-to-fine approaches. AnyTool [6] implemented a multi-level retrieval system based on the RapidAPI dataset, constructing separate retrievers for “category-tool-API” hierarchies. ToolRerank [32] leveraged pre-trained BERT models for semantic matching, while COLT [23] employed specialized language models for tool selection. Re-Invoke [5] introduced key information extraction from user queries before tool matching.
Figure 2: Illustrative example of MCP-Zero’s iterative proactive invocation. The model progressively identifies capability gaps and requests tools across three domains on-demand.
Limitations of Current Approaches. Despite reducing context overhead, existing retrieval-based methods face fundamental limitations. First, they usually rely on single-round matching based on initial user queries: if the retrieval fails, task completion becomes impossible. Second, and more critically, user requests are often broad and composite in nature (e.g., “write and run code to crawl AI repositories from GitHub”), requiring decomposition into multiple subtasks that cannot be solved by any single tool. Such scenarios demand progressive task breakdown where the model iteratively identifies capability gaps and requests appropriate tools on-demand, a challenge that current “query-once, retrieve-once” paradigms cannot adequately address.
2.3 Model Context Protocol
The Model Context Protocol (MCP) is an open standard introduced in 2024 to enable secure and uniform access to external tools and services for large language models [2]. By defining a standardized interface through JSON-RPC message exchange, MCP addresses the fragmentation problem in AI tool integration, where different platforms previously required custom connectors and proprietary protocols. This standardization has led to rapid adoption across major AI platforms and development environments, facilitating the creation of hundreds of MCP servers spanning diverse domains including file systems, databases, web services, and specialized APIs.
However, this expansion creates significant context overhead, as traditional approaches inject all server JSON-Schemas into system prompts simultaneously. Current solutions for MCP context optimization remain limited and follow conventional retrieval paradigms [7]. Given the enhanced capabilities of modern LLMs and MCP’s potential for agentic AI, we argue that existing methods fail to fully leverage contemporary language models’ tool-calling capabilities, motivating our MCP-Zero framework for dynamic, on-demand tool discovery.
3 Method: MCP-Zero
We introduce MCP-Zero through two key aspects: the comprehensive framework design (Section 3.1) and theoretical analysis of proactive retrieval (Section 3.2). Our approach fundamentally shifts from passive tool injection to active tool discovery, enabling LLMs to dynamically construct task-specific toolchains across diverse domains with minimal context overhead.
3.1 Proactive Tool Discovery Framework
MCP-Zero is a proactive agent framework that enables LLMs to dynamically construct task-specific toolchains through on-demand tool retrieval. The framework operates through three core components that work synergistically to address the context overhead and multi-domain coordination challenges of existing approaches.
Overall Workflow. Given a user query such as “Debug my code: src/train.py”, the LLM analyzes the task and autonomously determines when external tool assistance is needed. As shown in Figure 2, the model progressively breaks down the complex debugging task into subtasks: reading the file, analyzing and fixing the code, and validating the fix through execution. At each step, instead of relying on pre-injected tool schemas, the model proactively generates structured tool requests. These requests are processed through hierarchical vector routing to retrieve the most relevant tools, which are then injected into the context for immediate use. The model iteratively repeats this process throughout the conversation, constructing a cross-domain toolchain spanning filesystem operations, code editing, and command execution.
Proactive Tool Request. The foundation of our framework lies in returning tool requirement specification authority to the LLM itself. When the model identifies a capability gap that requires external assistance, it generates a structured request block:
<tool_assistant>
   server: ...  # Platform/permission domain
   tool: ...    # Operation type + target
</tool_assistant>
This mechanism enables the model to express tool needs spontaneously as they arise during task execution. Compared to raw user queries, the model-generated requests ensures better semantic alignment with tool documentation. The server field specifies the platform or permission domain requirements, while the tool field describes the desired operation type and target.
We design our framework around these two fields because they align naturally with the MCP specification, which mandates that all servers and tools provide descriptive documentation. This inherent requirement ensures consistent semantic information availability across the entire MCP ecosystem, making our retrieval approach universally applicable without additional metadata engineering.
Importantly, the model can generate multiple such requests throughout a single conversation, with each request triggering an independent retrieval process.
Hierarchical Vector Routing. To efficiently locate relevant tools from thousands of candidates, we employ a two-stage coarse-to-fine retrieval algorithm, using OpenAI text-embedding-3-large embeddings for semantic similarity matching.
The system first filters candidate MCP servers by matching the server field against server descriptions. Since server descriptions are typically brief single sentences, we construct extended summaries that include comprehensive usage examples essential for accurate matching (detailed in Section 4). We then perform matching against both the original descriptions and these enhanced summaries, taking the higher similarity score between the two approaches. This dual-matching strategy leverages both original MCP documentation and enhanced summaries to improve retrieval precision.
Subsequently, tools within selected servers are ranked based on semantic similarity between the tool field and tool descriptions. The final ranking score combines both server-level and tool-level similarities:
$\text{score}=(s_{\text{server}}\times s_{\text{tool}})\times\max(s_{\text{server}},s_{\text{tool}})$ (1)
where $s_{\text{server}}$ and $s_{\text{tool}}$ represent cosine similarities at server and tool levels respectively. This scoring mechanism ensures that high similarity in either dimension contributes significantly to the final ranking, enabling effective recall of highly relevant tools. The system returns the top-$k$ tools, with dynamic adjustment possible when similarity scores are closely clustered. In our experiments, we achieve high accuracy with top-$1$ retrieval, though we can configure larger $k$ values to enhance fault tolerance when needed.
Iterative Proactive Invocation. Unlike traditional single-round retrieval approaches, MCP-Zero supports iterative tool discovery throughout task execution. After receiving retrieved tools, the model autonomously evaluates their adequacy for the current subtask. If the returned tools are insufficient or inappropriate, the model can refine its request specification and reinitiate retrieval, providing natural fault tolerance and self-correction capabilities. This iterative process continues until the model determines that either (1) suitable tools have been found and the task can proceed, or (2) no appropriate tools exist and the task should rely on the model’s parametric knowledge.
The framework’s key advantage lies in its ability to construct cross-domain toolchains dynamically. For complex tasks requiring coordination across multiple domains (e.g., filesystem access, code generation, and command execution in Figure 2), the model can progressively identify and request tools from different servers as subtask requirements become clear, avoiding the context overhead of pre-loading comprehensive tool collections while maintaining high task completion accuracy. MCP-Zero represents a fundamental shift from “predefined toolset” to “dynamic on-demand tool discovery” for the community.
3.2 Theoretical Analysis
We provide a theoretical analysis of MCP-Zero’s advantages over traditional approaches through formal probability modeling and complexity analysis.
Problem Formulation. Let $T=\{t_{1},t_{2},...,t_{n}\}$ denote the complete tool collection, $q$ represent the user query, $c$ the conversation context, and $t^{*}$ the optimal tool selection. In traditional approaches, models must select tools from the entire collection simultaneously:
$P_{\text{traditional}}(t^{*}|q,c,T)=\frac{P(q,c|t^{*},T)P(t^{*}|T)}{\sum_{t_{i}\in T}P(q,c|t_{i},T)P(t_{i}|T)}.$ (2)
In contrast, MCP-Zero employs a two-stage process where models first generate structured requests $r$, then retrieve tools based on these requests:
$P_{\text{proposed}}(t^{*}|q,c)=\sum_{r}P(t^{*}|r)P(r|q,c),$ (3)
where $P(r|q,c)$ represents the probability of generating request $r$ given the query and context, and $P(t^{*}|r)$ denotes the probability of retrieving the optimal tool given the request.
Complexity Bottlenecks of Traditional Approaches. Traditional methods face three fundamental limitations that compound as tool collections scale:
Computational Complexity Growth: Models must evaluate all $n$ tools simultaneously, resulting in $O(n)$ computational complexity. This linear growth becomes prohibitive as tool ecosystems expand.
Attention Dilution Effects: Assuming finite attention resources, each tool receives approximately $\frac{1}{n}$ of the model’s attention. Introducing a noise factor $\eta(n)\propto\log(n)$ that increases with collection size, the effective selection probability degrades to:
$P_{\text{noise}}(t^{*}|q,c,T)=P_{\text{traditional}}(t^{*}|q,c,T)\cdot(1-\eta(n)).$ (4)
Semantic Distribution Mismatch: User queries and formal tool descriptions exhibit semantic distribution gaps quantifiable through KL divergence $D_{\text{KL}}(P_{q}\|P_{t})$, which reduces matching accuracy in vector-based retrieval systems.
Advantages of Proactive Request Mechanism. MCP-Zero addresses these limitations through two key mechanisms:
Search Space Compression: The hierarchical routing operates in two stages: first filtering among $m$ servers where $m\ll n$, then matching within the filtered tool subset $T_{\text{filtered}}$ from top-m selected servers. This reduces total complexity from $O(n)$ to $O(m+|T_{\text{filtered}}|)$ where $|T_{\text{filtered}}|\ll n$.
Semantic Consistency Enhancement: Model-generated requests $r$ exhibit higher semantic alignment with tool descriptions compared to raw user queries:
$D_{\text{KL}}(P_{r}\|P_{t})<D_{\text{KL}}(P_{q}\|P_{t}),$ (5)
which enhances vector matching accuracy.
Information-Theoretic Analysis and Probability Gains. From an information-theoretic perspective, traditional methods require decisions over high-entropy spaces. The decision entropy for traditional approaches is:
$H_{\text{traditional}}(T)=-\sum_{t_{i}\in T}P(t_{i})\log P(t_{i})\approx\log(n).$ (6)
MCP-Zero’s hierarchical filtering reduces the effective decision entropy to:
$H_{\text{proposed}}(T_{\text{filtered}})$ $=-\sum_{t_{i}\in T_{\text{filtered}}}P(t_{i})\log P(t_{i})$ (7)
$\approx\log(|T_{\text{filtered}}|)\ll\log(n).$
This entropy reduction translates to measurable probability gains. Defining the probability gain as $\Delta P=P_{\text{proposed}}(t^{*}|q,c)-P_{\text{traditional}}(t^{*}|q,c,T)$. In the traditional uniform, the baseline probability for selecting any specific tool approaches $\frac{1}{n}$. However, the attention dilution effect introduced by processing large tool collections degrades this baseline by the noise factor $\eta(n)$, yielding an effective success probability bounded by $\frac{1}{n}(1-\eta(n))$.
In contrast, MCP-Zero operates over a compressed search space of size $m+|T_{\text{filtered}}|$ where $m\ll n$ and $|T_{\text{filtered}}|\ll n$. More critically, the semantic alignment improvement quantified by the KL divergence reduction introduces a multiplicative enhancement factor $\alpha>1$:
$P_{\text{proposed}}(t^{*}|q,c)$ $\approx\frac{\alpha}{m+|T_{\text{filtered}}|}$ (8)
$P_{\text{traditional}}(t^{*}|q,c,T)$ $\approx\frac{1-\eta(n)}{n}$ (9)
Since $m+|T_{\text{filtered}}|\ll n$ and $\alpha>1$ due to improved semantic consistency, we have $\Delta P>0$ for practical tool collection sizes. Furthermore, as $n$ increases while $m$ and $|T_{\text{filtered}}|$ remain relatively stable (reflecting the hierarchical structure of tool ecosystems), the probability gain $\Delta P$ grows approximately linearly with $n$, demonstrating MCP-Zero’s superior scalability as tool ecosystems expand.
4 Dataset: MCP-Tools
To support our proactive tool discovery framework and enable comprehensive evaluation, we construct MCP-tools, the first retrieval-oriented dataset in the MCP domain. Unlike existing MCP evaluation frameworks such as MCPBench [13] that focus on server availability and latency testing, our dataset is specifically designed to facilitate semantic tool discovery and retrieval for large language models.
4.1 Data Collection and Scope
We systematically collected MCP server information from the official Model Context Protocol servers repository222github.com/modelcontextprotocol/servers (Tag: 2025.4.28, Commit: ad2d4e6). The repository contains three categories of servers: (1) 20 reference servers maintained by the MCP team, (2) 114 third-party official servers, and (3) 262 community-contributed servers, totaling 396 servers.
Our collection process involved extracting initial metadata including server names, brief descriptions, and repository URLs. Most servers link to GitHub repositories containing comprehensive documentation in README files, which include detailed server descriptions, tool specifications, usage examples, and implementation guidelines.
Figure 3: Needle-in-a-haystack test results demonstrating MCP-Zero’s performance under extreme scale conditions with varying tool collection sizes. The left column shows baseline methods with standard tool call schemas; the middle column shows our MCP-Zero approach; the right column shows MCP-Zero enhanced with one ICL example to guide the model’s description generation. Our method shows significant gains on Claude-3.5-Sonnet and Gemini-2.5-Flash, while GPT-4.1 shows no improvement due to its already strong baseline performance.
4.2 Data Processing and Quality Control
We implemented a rigorous filtering process to ensure data quality and completeness. Servers were excluded if they: (1) lacked MCP-compliant tool definitions, (2) missing documentation, or (3) failed to provide sufficient semantic information for retrieval purposes. This filtering process reduced our dataset from 396 to 308 high-quality servers.
For structured information extraction, we employed Qwen2.5-72B-Instruct [24] with manually crafted few-shot in-context learning examples to guide the extraction process. The extraction follows a standardized JSON schema:
  "server_name": string,
  "server_description": string,
  "server_summary": string,
  "tools": [
      "name": string,
      "description": string,
      "parameter": {
        "param1": "(type) description1",
        "param2": "(Optional, type) description2"
A critical component of the MCP-tools dataset is the server_summary field, which we generate by leveraging LLMs to synthesize comprehensive summaries from README content. These summaries capture server capabilities, features, and usage patterns while excluding operational details like installation instructions. This design enables effective matching regardless of whether the model expresses tool requirements in formal descriptions or more specific, invocation-oriented language, thereby improving retrieval precision across diverse query formulations.
4.3 Dataset Statistics and Characteristics
The final MCP-tools dataset comprises 308 servers and 2,797 tools, representing a comprehensive coverage of the MCP ecosystem. The tool distribution exhibits significant variance: the mean number of tools per server is 9.08 (median: 5.0, standard deviation: 11.40). Over half of the servers (162 servers) contain 5 or fewer tools, while some specialized servers include over 60 tools, reflecting the diverse nature of MCP applications ranging from simple utilities to comprehensive API suites.
To support efficient retrieval, we pre-compute embeddings for all textual descriptions and summaries using OpenAI text-embedding-3-large, creating a searchable vector index that enables rapid semantic matching during the hierarchical routing process. These pre-computed embeddings are included as part of the dataset release.
To our knowledge, this represents the first dataset specifically designed for semantic tool retrieval in the MCP domain, complementing existing evaluation frameworks like MCPBench [13] by focusing on discovery rather than performance testing. The dataset’s design directly supports our MCP-Zero framework while remaining general enough to benefit other retrieval-based tool selection approaches. By providing both original MCP documentation and enhanced semantic summaries, MCP-tools enables researchers to explore various retrieval strategies and evaluate their effectiveness across diverse tool domains.
Figure 4: Token efficiency comparison in needle-in-a-haystack experiments. The graph shows the average token cost per successful retrieval across different collection sizes.
5 Experiments
5.1 Existing Datasets and Their Limitations
Before presenting our experimental evaluation, we analyze existing tool-calling datasets and explain why they are insufficient for evaluating our MCP-Zero framework.
API-Bank [12] provides tool information with multi-turn conversations and human-annotated test sets, making it the most suitable existing dataset for our evaluation. ToolBench [20] collected 16,464 REST APIs from RapidAPI Hub but many APIs lack essential descriptions. ToolBank [15] improved upon ToolBench but relies on model-generated data. ToolAlpaca [28] synthesized 3,938 instances from 400 APIs but targets scenarios mismatched with contemporary use cases. The APIBench1 from Gorilla [18] contains 1,645 APIs but uses entirely GPT-synthesized conversations. Another APIBench2 [19] is proposed, but its application scenarios have low relevance to MCP tool calls.
Existing datasets have limitations for evaluating MCP-Zero: API scenarios differ from contemporary MCP use cases, lack hierarchical server-tool organization, or miss critical evaluation fields. Therefore, we use API-Bank as a reference and create the MCP-tools dataset for comprehensive evaluation. We are continuing to investigate other available datasets and conducting further validation.
Collection Method Claude-3.5 GPT-4.1 Gemini-2.5 Avg. Tokens $\downarrow$
Single-turn Conversation
Domain Q.Retrieval 71.63
Standard 97.60 98.08 92.79 312.4
MCP-Zero 96.15 96.62 97.12 111.0 (-64.47%)
Full Q.Retrieval 71.63
Standard 69.23 94.71 94.23 6308.2
MCP-Zero 95.19 95.19 96.63 111.0 (-98.24%)
Multi-turn Conversation
Domain Q.Retrieval 65.05
Standard 100.00 99.46 91.40 406.4
MCP-Zero 91.40 93.01 93.01 159.0 (-60.84%)
Full Q.Retrieval 65.05
Standard 60.22 93.01 92.47 6402.2
MCP-Zero 90.32 92.47 94.62 159.0 (-97.52%)
Table 1: APIBank evaluation results comparing MCP-Zero with standard tool calling methods across different scenarios. The accuracy is based on the top-$1$ result. Results show accuracy (%) for three LLMs and average token consumption. “Q.Retriaval” indicates retrieval based on user query.
5.2 Needle-in-a-Haystack Evaluation
To evaluate MCP-Zero’s ability to accurately retrieve tools from large-scale collections under extreme context conditions, we conduct needle-in-a-haystack experiments on our MCP-tools dataset.
Experimental Setup. We construct test scenarios by injecting 1 to 2,797 tools into the environment, selecting task descriptions from various positions as queries, and requiring models to retrieve the target tool. This setup simulates the challenging scenario where relevant tools are buried within massive tool collections. We compare three approaches:
  • Baseline: Standard tool call schemas with all tools injected into context
  • MCP-Zero: Our proactive retrieval approach
  • MCP-Zero + ICL: MCP-Zero enhanced with one in-context learning example to guide description generation
Results Analysis. As shown in Figure 3, our method demonstrates significant performance gains on Claude-3.5-Sonnet and Gemini-2.5-Flash, while GPT-4.1 shows no improvement due to its already strong baseline performance across all tool collection sizes. Figure 4 illustrates the dramatic difference in token consumption between traditional approaches and MCP-Zero. While standard tool call methods exhibit exponential growth in token costs as the number of MCP tools increases, MCP-Zero maintains consistently low token usage.
5.3 APIBank Evaluation
To validate MCP-Zero’s effectiveness in realistic conversational tool retrieval scenarios, we conduct comprehensive experiments on the APIBank dataset.
Experimental Setup. We extract description information from the APIBank level-1 dataset for retrieval tasks, get 48 unique tools in total, and process it for our evaluation framework. Since APIBank organizes data by individual APIs without server-level hierarchy, we directly retrieve tools without the server filtering stage.
We evaluate across two key dimensions: conversation context and tool collection scope. For conversation context, we test both single-turn scenarios (one user query for one response) and multi-turn scenarios (extended conversations with multiple exchanges). For tool collection scope, we examine both domain collections (using a curated subset of tools relevant to the specific domain) and full collections (retrieving from the complete set of all available tools).
Results Analysis. As shown in Table 1, we can conlude the following findings:
  • Extreme Context Efficiency. MCP-Zero cuts prompt length by 60–98% across all settings (e.g. 111 vs. 6.3k tokens in the full single-turn case), validating its ability to ”pay for tools only when they are needed”.
  • Robust Scalability. When moving from a hand-curated Domain subset to the Full tool pool (40x more APIs), standard schema-injection accuracy on Claude-3.5 plummets from 97.60 to 69.23 (single-turn) and 100.00 to 60.22 (multi-turn); MCP-Zero instead keeps accuracy at 95.19 / 90.32 respectively, demonstrating strong resilience to attention dilution.
  • Multi-turn Consistency. MCP-Zero maintains high accuracy over conversation rounds ($\leq$3drop from single- to multi-turn), whereas standard methods degrade sharply once the context accumulates previous calls and larger tool sets.
  • Necessity of Proactive Requests. Pure query-retrieval baselines stall at 65–72 % accuracy, confirming that letting the model author semantically aligned requests is crucial.
Experiments on APIBank corroborate our claims: MCP-Zero delivers near-optimal or superior tool-selection accuracy while slashing context usage by up to two orders of magnitude, remaining robust in both single- and multi-turn conversations and under massive tool-pool scaling. These results highlight proactive, iterative tool discovery as a practical path toward scalable, cost-efficient agent systems.
6 Discussion
In this section we reflect on how the MCP‐Zero paradigm can be adopted by other researchers (§6.1), analyse the surprisingly gain from a single in–context example (§6.2), and position MCP‐Zero with respect to the contemporaneous Alita system, outlining a promising path toward self-improving agentic AI (§6.3).
6.1 Cookbook: Integrate MCP-Zero Into Agent
MCP-Zero is fundamentally a simple yet effective approach that we hope will benefit the broader MCP community. The core methodology distills into three straightforward steps: prompting models to proactively request tools, maintaining a lightweight tool index with semantic descriptions, and leveraging the improved semantic alignment for high-precision retrieval. Below we provide a practical guide for integrating these ideas into existing agent frameworks.
Step 1 – Prompting the LLM to ask for tools. Give the model an explicit “permission” to declare missing capabilities. In practice this is a system instruction such as:
If the current task cannot be solved with your own knowledge, emit a <tool_assistant> block specifying the server domain and the tool operation you require.
In addition, the output structure needs to be specified as we mentioned in Section 3.1. This step aims to stimulate the model’s ability to ”proactively” propose requirements.
Step 2 – Curate a lightweight MCP-style tool index.
Firstly, choose a scope based on your needs: the entire MCP-tools collection, a vertical slice (e.g. databases only), or your in-house APIs. Then, for every server/tool:
  • extract the name and description from metadata;
  • optionally let a strong LLM generate an enhanced summary that emphasises capabilities and usage patterns;
  • store all texts in a vector store with pre-computed embeddings such as text-embedding-3-large.
Step 3 – Marry model output and retrieval.
When the agent emits a <tool_assistant> block:
  • Match the server field against server descriptions and summaries; take top-$m$ candidates.
  • Within each candidate server, rank tools by the tool field with the tool description embeddings.
  • Feed the best (or top-$k$) JSON-schemas back to the LLM.
Because the request text is already semantically aligned with the documents, retrieval precision is higher than “user query → API doc” matching, maintaining performances while significantly conserving context.
6.2 Why Does a Single ICL Example Help?
In §5.2 we observed that adding one in-context example (“ICL-1”) helps lifting needle-in-haystack accuracy marginally. We hypothesise two simple but potent effects:
  1. 1.
    Stylistic anchor. Our base prompt merely says “output the server and tool you need”, but gives no example of how the sentence should look like. The single in-context sample provides the writing style as the reference, helping the generated requests land much closer to the curated descriptions, thus semantic matching becomes easier.
  2. 2.
    Semantic grounding. The example also clarifies the meaning of each field, helping the model understand the specific definitions of MCP server and tool, thereby limiting its expression scope. After seeing this, the model reliably emits phrases such as filesystem_read instead of a vague “read the file”, sharply reducing semantic mismatch.
In short, a tiny demonstration patch acts as a schema anchor; future work could replace ICL with a short grammar-based decoder rule, but the one-shot approach is free and highly effective.
6.3 Synergy with Alita: Using and Making Tools
Concurrently, Alita [22] proposes a united manager agent that creates its own toolchain: it web-searches for code, clones GitHub repos, builds environments, and executes the resulting programs to accomplish tasks. We were pleasantly surprised by the contribution of this article, and found that the two lines of work are complementary:
  • MCP-Zero: efficiently finds and invokes existing tools
  • Alita : automatically builds missing tools on-the-fly
MCP-Zero and Alita address complementary halves of the same problem: the former maximises tool discovery while the latter maximises tool creation. When combined, they form a virtuous loop: an agent first proactively discover tools from all available resources; if none fits, it switches to Alita’s workflow to synthesize a new one, then registers the freshly built tool for the community. We believe such a pipeline is a compelling direction toward self-evolving, cost-aware agentic AI systems.
6.4 Future Work
While MCP-Zero demonstrates significant improvements in tool retrieval efficiency and accuracy, several promising directions warrant further investigation:
Enhanced Experimental Validation. Future work should expand evaluation across diverse domains. We plan to conduct comprehensive experiments on additional datasets to validate generalizability.
Advanced Matching Algorithms. The current semantic similarity approach could be enhanced. We envision incorporating multi-modal descriptions (e.g., code examples, usage patterns, parameter schemas) into the retrieval process, and exploring usage co-occurrence patterns for improved contextual understanding.
MCP Server Implementation. A natural extension involves packaging MCP-Zero as a dedicated MCP server providing tool discovery services. This ”meta-server” would expose standardized APIs for proactive tool retrieval, enabling seamless integration into existing MCP ecosystems and serving as a centralized discovery hub for distributed tool collections.
Multi-Agent Orchestration. MCP-Zero’s proactive discovery approach could enable better multi-agent collaboration. Future work could investigate how different agents can automatically discover and share tools with each other, allowing them to work together more effectively on complex tasks that require diverse capabilities.